OpenStudy (anonymous):

6, 12, 9, 18, 15 a. 45 c. 12 b. 23 d. 30

OpenStudy (anonymous):

make a conjecture about the next number in the sequence

OpenStudy (anonymous):

Well you can see that 6*2=12, and then 12-3=9 And you can see that 9*2=18, and 18-3=15 So, what is 15*2?

OpenStudy (anonymous):

30

OpenStudy (mathmate):

Hint: 6, 12, 9, 18, 15, ?

OpenStudy (anonymous):

Yup, so 30 would be your next number. :o) Good job!

OpenStudy (anonymous):

thanks! @Butterfly16